OpenStudy (anonymous):

The greatest common factor of 54x and 96xy is A. 13x B. 2x C. 9x D. 6x

OpenStudy (anonymous):

d

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D. 6x

OpenStudy (anonymous):

omg im soo smart :)

OpenStudy (anonymous):

and im soo bored ughh

OpenStudy (anonymous):

54 = 2*3*3*3 96 = 2*2*2*2*2*3 matching prime factors = 2 and 3 GCF = 6